# uDistil-Whisper: Label-Free Data Filtering for Knowledge Distillation via Large-Scale Pseudo Labelling

### v2

@@ 4.2 Baselines
We conduct evaluations on various multilingual and monolingual speech recognition models using different varieties of Arabic from the datasets mentioned earlier, encompassing both standard and accented MSA, as well as multiple Arabic dialects. We organize these systems into the following categories:
4.2.1 Supervised Models and Commerical Systems.
We evaluate two supervised baselines and a fine-tuned Whisper model for Arabic ASR. The baselines include Wav2Vec2-XLS-RConneau et al. (2020); Babu et al. (2021), trained on CV8.0, and HuBERTHsu et al. (2021), trained on MGB-3 Ali et al. (2017) and a 5.5-hour Egyptian Arabic speech corpus. We fine-tune the third model, whisper-large-v2, on CV11.0 and MGB-2. All models are evaluated using the test and dev splits of MGB-2 and CV11.0, as detailed in Section 4.1.
@@ 4.2.2 Zero-Shot Models
SeamlessM4T. SeamlessM4T models excel at generating high-quality transcripts across languages Communication et al. (2023b), but evaluations on them typically focus on English. To fill this gap, we test three of its versions – seamless-m4t-medium (SM4T-M), seamless-m4t-v1-large (SM4T-L-v1), and seamless-m4t-v2-large (SM4T-L-v2) – under a zero-shot setting for Arabic ASR, using the default parameters from the model’s inference pipeline.
@@ 4.2.3 Distilled Models
To compare our results with an equivalent supervised setup, we take two variants (in terms of size) of distilled models presented in Waheed et al. (2024). Apart from the reported results, we extend the evaluation on SADA’s test and valid splits. We take 16-16 and 32-16 models trained with a WER threshold of 80 on each of 100K and 500K segments. Waheed et al. (2024) find 80 as the optimal threshold which gives the best results. We also take models trained without any filtering and treat those as a lower bound.
@@ 4.3 Experimental Setup
During evaluations, we adhere to the default decoding parameters specified by the original model implementations unless noted otherwise. Throughout our experiments, we maintain a maximum sequence length of 225 and use WER and CER as the primary metrics for assessing performance.
For distillation, we select roughly 27% of the top segments based on the metrics, to duplicate the amount used in a supervised setup when $\lambda=80\%$. We refrain from conducting hyperparameter search due to computational constraints, and instead apply the configuration detailed in Gandhi et al. (2023b). The distillation process and its key parameters are detailed in Table 5 in Appendix 1.2.
@@ 5 Results and Discussion
Table 2: WER/CER scores after normalization and removing diacritics. All baseline distilled models (dw-) are trained with a filtering threshold of 80 if not specified. We report the score on the test split of each dataset. Abbreviations. W - Whisper, FT - Finetuned, M - Medium, L - Large, S - Small, U - Unsupervised, D - Distil.
@@
We report the baseline results from Waheed et al. (2024) and evaluate all models on SADA’s test and validation splits. We do the distillation with two variants of student models. The first student has 16 of each of the encoder and decoder blocks (UDW-16-16) and the second one has 32 encoder and 16 decoder blocks (UDW-32-16). Following Waheed et al. (2024), we experiment with 100K and 500K segments. We report our main results in Table 2, and the averages across different evaluation setups in Table 3. Results on the top five varieties [^3] We refer dialect labels in SADA as varieties as not all labels are necessarily dialects. of SADA are provided in Table 4. The reported results are computed on normalized text that includes removing diacritics. Results on the orthographic text (no normalization) can be found in Appendix 1.2.
Waheed et al. (2024) show that their supervised models perform well on in-distribution evaluation, however, their performance degrades considerably under novel and challenging conditions. Our evaluation of these supervised models shows similar results. We find HuBERT trained on MGB3 Ali et al. (2017) corpus outperforms all other supervised baselines across different evaluation setups.
@@
In zero-shot models, we compare our results with MMS Pratap et al. (2023a), three variants of SeamlessM4T family models (medium, large-v1, large-v2), and four Whisper family models. We extend the evaluation to SADA data for these models as well. While larger variants of both SeamlessM4T and Whisper perform on standard benchmark setups like Common Voice and Fleurs they perform poorly on the novel and challenging settings. For example, the best zero-shot model SeamlessM4T-large-v2 gives just 7.6% WER on Fleurs compared to 65.0% WER on the SADA test set as shown in Table 2. This underscores the inadequate evaluation and limited generalization capability of zero-shot models in unseen and challenging settings.
@@
Further, we take the reported results of distilled models in Waheed et al. (2024) and extend its evaluation to SADA data. As demonstrated in the reported results, the best-distilled model DW-32-16++ outperforms all other models including its teacher (large-v2), best of Whisper (large-v3), and best of baselines (SeamlessM4T-large-v2).
@@
We trained two versions of teacher models using datasets of 100K and 500K segments, respectively. Approximately 27% of the examples were removed from each dataset based on specific metrics, aligning our data set up with the one used in Waheed et al. (2024), which employs a filtering threshold of 80. Consequently, we consider models trained with this threshold of 80 to represent an upper bound, while models trained on the original, unfiltered pseudo labels serve as our baseline for comparison.
@@
From our experiments, we observed that two of our best measures, sonar-sim and proxy-ref based filtering, perform comparably to or even surpass the equivalent supervised setups. For instance, the model DW-32-16 trained on 100K examples with a filtering threshold of 80 achieved a WER of 35.3% on the benchmark test split, 66.6% on the SADA test split, and 61.1% averaged om five novel dialects. By comparison, our best-performing model in a similar configuration, UDW-32-16-sonar, achieves 35.5% WER on the benchmark test split, 58.7% on the SADA test, and 60.6% on five dialects. This reflects an almost 3% improvement in average score compared to the supervised setup (51.6 vs 54.3) and more than a 7% improvement over the no-filter setup, as detailed in Table 3.
@@
Additionally, our experiments show that the smaller student model (UDW-16-16) performs better with the proxy-ref as a filtering measure. More specifically, the UDW-16-16, experimented with 100K examples and utilizing proxy-ref to filter out low-quality pseudo labels, matches the results of supervised setups (58.5% vs 57.6%) and shows over a 6% improvement compared to setups where pseudo labels are not filtered (64.2%).
@@
Table 3: Average WER/CER across different evaluation datasets. WER/CER scores after normalization and removing diacritics.
@@
Following Waheed et al. (2024), we also expand our dataset from 100K to 500K segments, equivalent to approximately 500 hours of data. As with our other experiments, we filter out approximately 27% of low-quality examples from our training data to align with the quantity used in supervised baselines. We train two models, UDW-16-16 and UDW-32-16, employing proxy-ref and sonar-sim as filtering criteria.
As shown in Table 3, our UDW-32-16++ model, trained with 500K segments, performs comparably with the supervised counterpart DW-32-16++. For instance, UDW-32-16++ with proxy reference filter achieves WER of 51.4%, whereas DW-32-16++ achieves 50.0% WER. Notably, our smaller student model on 500K segments (UDW-16-16++) results in superior performance compared to models trained under a supervised filtering setup. For example, UDW-16-16++ with sonar-sim gives 53.9% WER compared to 56.2% WER for DW-16-16++. These results showcase the efficacy of our data filtering methods.
@@
Table 4: Results on top five dialects/categories in SADA data. WER/CER scores after normalization and removing diacritics.
We evaluate our models under various unseen conditions, including five novel dialects that our models have not encountered before. Additionally, we compute category-wise error rates for both the SADA test and valid splits. We present results on top five categories  [^4] We refer fategories as dialectal and non-dialectal labels in SADA data. As detailed in Table 4, our best model consistently outperforms all others when averaged across these novel dialects.
Furthermore, our top-performing model, UDW-32-16++, demonstrates superior performance over DW-32-16++ in the top five SADA categories. For instance, using proxy reference as a filtering measure, UDW-32-16++ achieves 58.06%, compared to DW-32-16++’s 59.42% WER averaged across top five categories in SADA test split. This highlights our ability to distill smaller models from large Whisper models, maintaining or improving performance while reducing size, all without requiring labeled data.
@@
We investigate the effectiveness of several of these metrics for filtering low-quality pseudo-labels in speech recognition tasks, specifically targeting instances with a WER higher than 80%, 40%, and 20%. To assess their efficacy, we calculate the area under the curve (AUC) as we show in Figure LABEL:fig:side-by-side for detecting low-quality examples. The results indicate that sonar-sim achieves an AUC of 0.77 for detecting examples with $WER>80$, demonstrating reasonably high discriminative power in identifying low-quality labels. The proxy-ref metric shows slightly better performance, with an AUC of 0.82, indicating robust capability in distinguishing between high and low-quality pseudo-labels. In contrast, the confidence-based measure yielded an AUC of 0.68, which shows relatively less discriminative power in filtering out undesirable labels. These findings highlight sonar embeddings and the proxy reference-based measure as promising tools for improving the quality of pseudo-labels in scenarios where ground truth data is unavailable.
@@ 6 Conclusion
In this study, we explore methods for distilling large Whisper models into smaller, more efficient ones. Our proposed techniques for filtering out low-quality examples eliminate the need for labeled data, bridging a gap in prior research and facilitating the creation of compact and effective speech recognition models for settings with limited labeled data. We conducted a comprehensive evaluation of our trained models, comparing them against various strong zero-shot and distilled models. Our results demonstrate that our models can outperform both the teacher model and models distilled with supervised data filtering setups.
We evaluate our models across a diverse range of Arabic varieties and dialects, assessing their generalization and adaptability to linguistic diversity. From large multilingual speech models (Whisper), we distill language-specific models optimized for Arabic ASR. Our evaluation spans three different setups, encompassing a wide array of linguistic varieties (ie: dialects), and reveals that despite their reduced size (25-50%), these distilled models consistently outperform their base counterparts and can compete with state-of-the-art models twice their size—all achieved without the need for labeled data.
@@ 7 Limitations
In this study, we distill small Whisper models from relatively large ones via pseudo-labeling and unsupervised data filtering. Our distilled models are computationally efficient and maintain a performance similar to or better than the base teacher model and models trained in a supervised data filtering setup. Unlike Waheed et al. (2024); Gandhi et al. (2023a), our approach does not utilize any labeled data in the distillation process, making it directly applicable in data-scarce settings. However, despite these advantages, we acknowledge several limitations in our work, which we outline below.
@@
Our main approach relies heavily on a robust reference model to serve as a proxy for filtering lower-quality pseudo labels. Specifically, we utilize SeamlessM4T-large-v2, a state-of-the-art model with 2.3 billion parameters, to generate proxy references which is then used to filter out low-quality data points. For similarity-based measures, we use SONAR to generate multimodal embeddings from speech and pseudo labels. These embeddings provide contextual similarity which is then utilized to discard low-quality pseudo labels. We use AceGPT (7B), to compute the log-likelihood of the pseudo labels is which leveraged to filter out low-quality examples.
Although with these measures we attain performance on par or better than the supervised setup, it’s important to highlight that each of these methodologies entails additional computational overhead.
@@
Evaluation. Arabic is a linguistically rich and complex language with over 400 million speakers Abdul-Mageed et al. (2021), resulting in its wide range of varieties and dialects. We evaluate all the models on ten different datasets representing different varieties, including five novel dialects collected and curated by native speakers and never seen before by any models. However, our varieties do not cover all Arabic-speaking regions. We aim to address this in future work by covering more varieties and dialects.
Distillation Training Data. We distilled four variants of student models using 100K and 500K segments of which approximately 25% are filtered. We see improvement going from 100K ($\approx$100100100100 hours) to 500K ($\approx$500500500500 hours) segments. As Gandhi et al. (2023b) shows going over 1000 hours results in a better model, we aim to study how distillation can be done under a low resource setting which is why we do not scale the data. Additionally, we also keep the WER threshold high (80) so that we remain close to a setting where no labeled data is available (even for filtering). It would be interesting, however, to see how distilled models may perform on unfiltered data in low-resource setting.

### v3

@@ 4.2 Baselines
We evaluate several multilingual and monolingual speech recognition models on different Arabic varieties, including standard and accented MSA, as well as various dialects. The models are grouped as follows:
4.2.1 Supervised Models and Commercial Systems.
We evaluate three baselines: Wav2Vec2-XLS-R Conneau et al. (2020); Babu et al. (2021) trained on CV8.0, HuBERT Hsu et al. (2021) trained on MGB-3 Ali et al. (2017) and a 5.5-hour Egyptian Arabic corpus, and a fine-tuned Whisper-large-v2 model trained on CV11.0 and MGB-2. These models have been used off the shelf and been trained by others.
@@ 4.2.2 Zero-Shot Models
SeamlessM4T. SeamlessM4T models excel at generating high-quality transcripts across languages Communication et al. (2023), but evaluations on them typically focus on English. To bridge this gap, we test three of its versions – seamless-m4t-medium (SM4T-M), seamless-m4t-v1-large (SM4T-L-v1), and seamless-m4t-v2-large (SM4T-L-v2) – under a zero-shot setting on Arabic ASR, using the default parameters from the model’s inference pipeline.
@@ 4.2.3 Distilled Models
To compare our results with an equivalent supervised setup, we evaluate two distilled model variants of different sizes. We extend the evaluation to SADA’s test and valid splits, using 16-16 and 32-16 models trained with a WER threshold of 80% on both 100K and 500K segments. Additionally, we include models trained without filtering as a lower bound. For fair comparison with supervised data filtering, we follow an experimental setup identical to prior work Waheed et al. (2024).
@@ 4.3 Experimental Setup
During evaluations, we adhere to the default decoding parameters specified by the original model implementations unless specified otherwise. Throughout our experiments, we maintain a maximum sequence length of 225 and use WER and CER as the primary metrics for quantitative evaluation of the models.
For distillation, we select roughly 27% of the top segments based on the metrics, to match the amount of data used in a supervised setup when $\lambda=80\%$. We do not conduct hyperparameter search due to computational constraints, and instead apply the configuration detailed in Gandhi et al. (2023). The distillation process and its key parameters are detailed in Table 6 in Appendix C.
@@ 5 Results and Discussion
Table 1: WER/CER scores after normalization and removing diacritics. All baseline distilled models (dw-) are trained with a filtering threshold of 80 if not specified. Best results are shown in bold. Second best results are underlined. We report the score on the test split of each dataset. Abbreviations. W - Whisper, FT - Finetuned, M - Medium, L - Large, S - Small, U - Unsupervised, D - Distil.
@@
We report the baseline results from Waheed et al. (2024) and evaluate all models on SADA’s test and validation splits. We distil two variants of student models. The first student has 16 of each of the encoder and decoder blocks (UDW-16-16) and the second one has 32 encoder and 16 decoder blocks (UDW-32-16). We experiment with 100K and 500K segments. We report our main results in Table 1, and the average across different evaluation setups in Table 2. Results on the top five varieties [^4] We refer labels in SADA as varieties as not all labels are necessarily dialects. of SADA are provided in Table 3. The reported results are computed on normalized text that includes removing diacritics. Results on the text with no normalization can be found in Appendix 3.2.
@@
In zero-shot models, we compare our results with MMS Pratap et al. (2023), three variants of SeamlessM4T family models (medium, large-v1, large-v2), and four Whisper models. We extend the evaluation to the SADA data for these models as well. While larger variants of both SeamlessM4T and Whisper perform well on standard benchmark setups like Common Voice and Fleurs, they perform poorly on novel and challenging settings, such as dialectal speech. For example, the best zero-shot model (SeamlessM4T-large-v2) has a 7.6% WER on Fleurs compared to 65.0% on the SADA test set as shown in Table 1. This underscores the inadequate evaluation and limited generalization capability of zero-shot models in unseen and challenging settings.
@@
As demonstrated in the reported results, the best-distilled model DW-32-16++ outperforms all other models including its teacher (large-v2), the best Whisper (large-v3), and the best baseline (SeamlessM4T-large-v2).
@@
We train two versions of teacher models using datasets of size 100K and 500K segments, respectively. Approximately 27% of the examples were removed from each dataset based on specific filtering metrics, matching the amount of data used in the supervised setup when the filtering threshold is 80.
@@
From our experiments, we observe that two of our best filtering measures, sonar-sim and proxy-ref, perform similar to or surpass the equivalent supervised setups. For instance, DW-32-16 trained on 100K examples with a filtering threshold of 80 achieves a WER of 35.3% on the benchmark test split, 66.6% on the SADA test split, and 61.1% averaged on five novel dialects. By comparison, our best-performing model in a similar configuration, UDW-32-16-sonar, achieves 35.5% WER on the benchmark test split, 58.7% on the SADA test, and 60.6% on five dialects. This reflects close to 3% improvement in average score compared to the supervised setup (51.6 vs 54.3) and more than a 7% improvement over the no-filter setup, as detailed in Table 2.
@@
In addition to that, our experiments show that the smaller student model (UDW-16-16) performs better with proxy-ref as a filtering measure. More specifically, the UDW-16-16, experimented with 100K examples and utilizing proxy-ref to filter out low-quality pseudo labels, and matches the results of supervised setups (58.5% vs 57.6%) and shows over 6% improvement compared to setups where pseudo-labels are not filtered (64.2%).
@@
Table 2: Average WER/CER across different evaluation datasets. Best results are shown in bold. Second best results are underlined. WER/CER scores after normalization and removing diacritics.
@@
We scale our dataset from 100K to 500K segments and we train two models, UDW-16-16 and UDW-32-16, employing proxy-ref and sonar-sim as filtering criteria. As shown in Table 2, our UDW-32-16++ model, trained with 500K segments, performs comparably with its supervised counterpart DW-32-16++. For instance, UDW-32-16++ with proxy-ref filtering achieves a WER of 51.4%, similar to DW-32-16++’s 50.0%. Notably, our smaller student model (UDW-16-16++) trained on 500K segments results in superior performance compared to models trained under supervised filtering. For example, UDW-16-16++ with sonar-sim gives 53.9% WER compared to 56.2 for DW-16-16++. These results showcase the efficacy of our data filtering methods.
@@
Table 3: Results on top five dialects/categories in SADA data. Best results are shown in bold. Second best results are underlined. WER/CER scores after normalization and removing diacritics.
We evaluate our models under various unseen conditions, including five novel dialects that our models have not encountered before. Additionally, we compute category-wise  [^5] By categories we refer to dialectal and non-dialectal labels in the SADA data. error rates for both the SADA test and valid splits. We present results on the top five categories. As detailed in Table 3, our best model UDW-32-16++ consistently outperforms all others when averaged across these novel dialects.
Furthermore, UDW-32-16++, demonstrates superior performance over DW-32-16++ in the top five SADA categories. For instance, when using proxy-ref as a filtering measure, UDW-32-16++ achieves 58.06% WER, compared to DW-32-16++’s 59.42% averaged across top (with most utterances) five categories in the SADA test split. This demonstrates our ability to (1) distill smaller models from larger Whisper models, (2) maintain or improve performance, and (3) reduce model size without relying on labeled data.
@@
We investigate the effectiveness of two of our best metrics for filtering low-quality pseudo-labels, specifically targeting instances with a WER higher than 80%, 40%, and 20%. To assess their efficacy, we calculate the area under the curve (AUC) (as shown in Figure LABEL:fig:side-by-side) for detecting low-quality examples. The results indicate that sonar-sim achieves an AUC of 0.77 for detecting examples with a WER $>80$, demonstrating reasonably high discriminative power in identifying low-quality labels. The proxy-ref metric shows a slightly better performance, with an AUC of 0.82, indicating robust capability in distinguishing between high and low-quality pseudo-labels. In contrast, the confidence-based measure yielded an AUC of 0.68, which falls behind the other measures’ discriminative power. These findings highlight sonar embeddings and the proxy reference-based measure as promising tools for improving the quality of pseudo-labels in scenarios where ground truth data is unavailable.
5.1 Experiments on Other Language.
Evaluation Dataset W-L-v2 DW-16-16 DW-32-16 UDW-16-16-proxy UDW-32-16-proxy
IID OpenBible 101.33/44.43 59.14/13.95 58.78/13.75 59.21/13.96 58.86/14.13
CommonVoice17 117.13/60.13 82.91/34.96 69.8/24.77 75.58/29.18 70.43/25.36
ALFAA 217.11/143.19 78.24/28.23 74.4/25.72 76.82/27.2 73.81/26.53
OOD DVoice 214.6/144.57 124.35/74.1 110.19/62.61 110.68/62.4 114.9/69.14
AMMI-LigAikuma 46.71/12.96 60.1/18.02 51.77/14.37 60.44/18.47 52.19/14.42
Fleurs 54.61/14.81 60.85/18.85 51.58/14.8 58.87/18.45 51.82/14.88
Table 4: WER/CER comparison on Swahili. Best results are shown in bold. Second best results are underlined.
To further validate the effectiveness of our approach, we conduct experiments on Swahili, a low-resource language. We collect over 100 hours of labeled speech data from a variety of sources, namely OpenBible Meyer et al. (2022), CommonVoice (Swahili subset) Ardila et al. (2020), ALFAA [^6] https://github.com/besacier/ALFFA_PUBLIC/tree/master/ASR/SWAHILI, DVoice Gauthier et al. (2016), AMMI-LigAikum [^7] https://github.com/besacier/AMMIcourse, and FLEURS (Swahili subset) Conneau et al. (2022).
We distill two models, UDW-16-16 and UDW-32-16, using our best filtering method: proxy-ref. The training data includes the train splits of OpenBible, CommonVoice, and ALFAA, and we evaluate the models on their respective test splits. We also test the models on three out-of-distribution (OOD) datasets: DVoice, AMMI-LigAikum, and FLEURS, which were not included in the training data.
We compare our distilled models to the teacher model and evaluate the performance of our unsupervised approach. The results show that our unsupervised distillation models perform on par with, or better than the supervised setup. Additionally, our distilled models outperform the teacher model by a significant margin on both IID and OOD datasets, demonstrating the utility of our approach in extremely low-resource settings. Specifically, the UDW-32-16 model achieves a WER/CER of 58.86/14.13% on the IID OpenBible dataset, compared to the teacher model’s 101.33/44.43%. On the OOD dataset FLEURS, UDW-32-16 attains a WER/CER of 51.82/14.88, significantly outperforming the teacher model’s 54.61/14.81. Across various datasets, our distilled models consistently outperform the teacher, with UDW-32-16 showing the best results overall. Table 4 presents the WER and CER scores for the different models and datasets.
These findings highlight the strength of our unsupervised data filtering approach, particularly in low-resource scenarios, where labeled data is scarce but the distilled models still perform robustly.
@@ 6 Conclusion
In this study, we explore methods for distilling large Whisper models into smaller, more efficient ones. Our proposed techniques for filtering out low-quality examples eliminate the need for labeled data, bridging a gap in prior research and facilitating the creation of compact and effective speech recognition models for settings with limited labeled data. We conduct a comprehensive evaluation of our trained models, comparing them against various strong zero-shot and distilled models. Our results demonstrate that our models can outperform both the teacher model and models distilled with supervised data filtering setups.
We evaluate our models across a diverse range of Arabic varieties and dialects, assessing their generalization and adaptability to linguistic diversity. From large multilingual speech models (Whisper), we distill language-specific models optimized for Arabic ASR. Our evaluation spans three different setups, encompassing a wide array of linguistic varieties (ie: dialects), and reveals that despite their reduced size (25-50%), these distilled models consistently outperform their base counterparts and can compete with state-of-the-art models twice their size—all achieved without the need for labeled data. We also apply our unsupervised distillation approach on Swahili, another low-resource language, which shows a similar performance and proves its efficacy in different setting.
@@ 7 Limitations
In this study, we distill small Whisper models from relatively large ones via pseudo-labeling and unsupervised data filtering. Our distilled models are computationally efficient and maintain a performance similar to or better than the base teacher model and models trained in a supervised data filtering setup. Unlike Waheed et al. (2024); Gandhi et al. (2023), our approach does not utilize any labeled data in the distillation process, making it directly applicable in data-scarce settings. However, despite these advantages, we acknowledge several limitations in our work, which we outline below.
@@
Our main approach relies heavily on a robust reference model to serve as a proxy for filtering lower-quality pseudo labels. Specifically, we utilize SeamlessM4T-large-v2, a state-of-the-art model with 2.3 billion parameters, to generate proxy references which is then used to filter out low-quality data points. For similarity-based measures, we use SONAR to generate multimodal embeddings from speech and pseudo labels. These embeddings provide contextual similarity which is then utilized to discard low-quality pseudo labels. We use AceGPT (7B), to compute the log-likelihood of the pseudo labels which is leveraged to filter out low-quality examples.
Although these measures allow attaining a performance on par or better than the supervised setup, it’s important to highlight that each of these methodologies entails additional computational overhead.
@@
Evaluation. Arabic is a linguistically rich and complex language with over 400 million speakers Abdul-Mageed et al. (2021), resulting in its wide range of varieties and dialects. We evaluate all the models on eleven different datasets representing different varieties, including five novel dialects collected and curated by native speakers and never seen before by any models. However, our varieties do not cover all Arabic-speaking regions. We aim to address this in future work by covering more varieties and dialects.
Distillation Training Data. We distilled four variants of student models using 100K and 500K segments of which approximately 25% are filtered. We see improvement going from 100K ($\approx$100100100100 hours) to 500K ($\approx$500500500500 hours) segments. As Gandhi et al. (2023) shows, going over 1000 hours results in a better model, we aim to study how distillation can be done under a low resource setting which is why we do not scale the data. Additionally, we also keep the WER threshold high (80) so that we remain close to a setting where no labeled data is available (even for filtering). It would be interesting, however, to see how distilled models may perform on unfiltered data in low-resource setting.